OpenStudy (anonymous):

use Lagrange Multipliers to solve. A rectangular tank with a bottom and sides but no top is to have volume 500 cubic feet. Determine the dimensions (length, width, height) with the smallest possible surface area.

OpenStudy (anonymous):

you need amistre for this. he is a lagrange multiplier user

OpenStudy (anonymous):

bookmarking for future reference.

OpenStudy (sriram):

l*b*h=500 l*b+2(b*h+l*h)=surface area diiferentate the first eq to get b*h dl +l*b dh + l*h db=0 differentiate the second to get